P12.4B Exercise

P12.4B (LO 34) AP Terry Lam and Chris Tan have a partnership agreement with the following provisions for sharing profit or loss:

  1. A salary allowance of $20,000 to Lam and $30,000 to Tan

  2. An interest allowance of 5% on capital balances at the beginning of the year

  3. The remainder to be divided between Lam and Tan on a 3:4 basis

The capital balances on February 1, 2020, for T. Lam and C. Tan were $100,000 and $120,000, respectively. For the year ended January 31, 2021, the partnership reported a loss of $30,000. The partnership also reported the following: T. Lam drawings of $12,000 and C. Tan drawings of $14,400.

Instructions

a. Prepare a schedule to show how the profit or loss is allocated to the two partners.

b. Prepare a statement of partners' equity for the year.

c. Prepare closing entries on January 31, 2021, to allocate the partnership profit and to close the drawings accounts.
Answer #1

Answer

Partners : Terry Lam & Chris Tan

The question states that the year ending result of the firm is Net Loss of $30,000. So, it is assumed that this loss is after meeting interest on capital and salary allowance.

Residuary profit sharing ratio is 3:4. That means, loss of $30,000*3/7 = $12,857 will be deducted from the net income of Terry Lam and the balance loss of $17,143.

Statement of Net Income to Partners

Particulars Terry Lam Chriss Tan Total
Salary allowance 20,000 30,000 50,000
Interest on Capital (5% of beginning Capital Account Balance) 5,000 6,000 11,000
25,000 36,000 61,000
Less : Loss for the year (12,857) (17,143) (30,000)
Net Income 12,143 18,857 31,000

Interest on capital

To Terry Lam = $100,000 * 5% = $5,000

To Chris Tan = $120,000 * 5% = $6,000

Since the date of withdrawal is not given, interest on opening capital is allowed for the whole year.

Statement of Partners' Equity for the year ended Jan 31, 2021

Particulars Terry Lam Chriss Tan Total
Capital Balance on Feb 1, 2020 100,000 120,000 220,000
Add : Additinal capital Introduced 0 0 0
Add : Net Income 12,143 18,857 31,000
112,143 138,857 251,000
Less : Withdrawals during the year (12,000) (14,400) (26,400)
Capital Balance on Jan 31, 2021 100,143 124,457 224,600

Closing Entries on Jan 31,2021

Date General Journal Debit Credit
31 Jan 2021 Terry Lam Capital Account 12,000
Chriss Tan Capital Account 14,400
Drawings Account 26,400
(Being Withdrawal adjusted against Capital Accounts of partners)
31 Jan 2021 Terry Lam Capital Account 12,857
Chriss Tan Capital Account 17,143
Profit & Loss Account 30,000
(Being net loss for the year transferred to the capial account of partners in the sharing ratio of 3;4)